OpenStudy (anonymous):

A downward force of 8.4N is exerted on a -8.8uC charge. What are the magnitude and direction of the electric field?

OpenStudy (anonymous):

\[E=F/q\] E= 8.4N/ -8.8 X 10^6 C= Upward

OpenStudy (anonymous):

what would be the unit?

OpenStudy (anonymous):

thank you very much

OpenStudy (prakharluv):

its unit will be Newton per coulomb

OpenStudy (prakharluv):

E=F/q E=8.4 N/-8.8*10^6=-9.5*10^-7 N/C -ve sign indicates the direction of Electric field is vertically upward.

OpenStudy (anonymous):

if the sign is positive the direction will be downward?

OpenStudy (prakharluv):

yup

OpenStudy (anonymous):

okay now i finally got the problem thank you. but this one is different: An electron is released from the rest in a uniform electric field and accelerates to the north at the rate of 115m/s^2. What are the magnitude and direction of electric field (E)?

OpenStudy (prakharluv):

we know that F=ma, so on putting the value of m and a you will get force... and we know that the charge on electron...simply put all the values in eqn. E=F/q=ma/q you will get ur ans

OpenStudy (anonymous):

the acceleration would be_115m/s^2 right? what would be the mass? and our q would be the electrons charge?

OpenStudy (prakharluv):

mass of electron is 9.1*10^-31kg and charge on electron is -1.6*10^-19C

OpenStudy (prakharluv):

ans will be -6.5*10^-48 N /C

OpenStudy (anonymous):

this: (115m/s^2) ( 9.1*10^-31kg) divided by -1.6*10^-19C ?

OpenStudy (prakharluv):

yeah

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

how did you get this ( 9.1*10^-31kg) ?

OpenStudy (prakharluv):

it is calculated value..you may learn it. u may also learn the charges and masses of another sub atomic particles

OpenStudy (anonymous):

is it constant? can u show me how did u get it?

OpenStudy (prakharluv):

yes it is constant... and for further information please concern with wikipedia here u will get ur ans...

OpenStudy (anonymous):

how did u get the exponent -48? i got -10 ?

OpenStudy (anonymous):

yes, -10

OpenStudy (anonymous):

how about this: What are the magnitude and direction of the electric force on an electron in a uniform electric field of strength 2360N/C that points due East. the charge of electron is 1.6 x10^(-19) ? how come? Force=electric field x charge electric field=2360

OpenStudy (prakharluv):

what are u asking alfers101?

OpenStudy (anonymous):

The charge of an electron is always the same, it is a given ( and it should be negative though). You have to plug in. Pay attention to the units, write the units of the given. You have to still determine the direction. F= qE; plug in ; and since it is negative it should be in the opposite direction.